OpenStudy (anonymous):

There are 6 dogs and 5 cats. In how many different orders can these animals be placed in line under the following conditions? i) Any animal can be next to any other animal. ii) The dogs and cats are lined up alternately. iii) The first and last animal in line must be a cat.

OpenStudy (anonymous):

_ _ _ _ _ _ _ _ _ _ _

OpenStudy (anonymous):

11 slots

OpenStudy (anonymous):

multiply

ganeshie8 (ganeshie8):

lets assume each dog is different and each cat is different

ganeshie8 (ganeshie8):

you can choose either a cat or dog in the first slot : 11 ways

ganeshie8 (ganeshie8):

from the remaining 10 catdogs, you can choose either a cat or dog in the second slot : 10 ways

ganeshie8 (ganeshie8):

...

OpenStudy (anonymous):

I thought only a cat?

OpenStudy (anonymous):

10.9.8.7.6.5.4.3.2.1

ganeshie8 (ganeshie8):

total number of ways of arranging these = 11*10*... 2*1 = 11!

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

please ganeshie8?

OpenStudy (anonymous):

Well ganeshie gave you the answer for the first one, it's 11!

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2. The row has two options, so 2*6!*5!

OpenStudy (anonymous):

not according to my choices

OpenStudy (anonymous):

i) 39,916,800, ii) 86,400, iii) 7,257,600 i) 42,068,100, ii) 88,700, iii) 7,406,800 i) 37,289,400, ii) 90,200, iii) 7127,400 i) 40,157,600, ii) 83,400, iii) 7,348,500

OpenStudy (anonymous):

Do you know what a factorial is?

OpenStudy (anonymous):

non-negative integer

OpenStudy (anonymous):

11! = 11*10*9*8*7...up to 1.

OpenStudy (anonymous):

36,288,000

OpenStudy (anonymous):

?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

Check again

OpenStudy (anonymous):

all those numbers

OpenStudy (anonymous):

On your calculator you should have a prb or probability button, and if you click that you will be able to find a exclamation mark "!" if you use that, it will save you time.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=11%21

OpenStudy (anonymous):

ok 39,916,800?

OpenStudy (anonymous):

so therefore answer has to be A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool :) thanks

OpenStudy (anonymous):

Np :P